OpenStudy (anonymous):

solve using the addition method 2x-y=8 4x+3y=36

OpenStudy (anonymous):

6x - 3y = 24 4x+3y=36 10x = 60 x = 6 y = 2x - 8 y = 4 x = 6 y = 4